The area of a rectangle is 14 square meters. Find the length and width of the rectangle if its length is 5 meters greater than its width.
Answer by mananth(16946)

width =x
length = x+5
Area = L*W
Area = x(x+5)
x(x+5)=14
x^2+5x=14
x^2+5x-14-0
x^2+7x-2x-14=0
x(x+_7)-2(x+7)=0
(x-2)(x+7)=0
x= 2 the positive value
width = 2 m
length = 2+5=7 m